OpenStudy (anonymous):

Finding gradients and first partials for the following: 1.f(x,y,z) = xe^z + ycos(x) 2. f(x,y,z) = (x+y+z)^8 3. f(x,y) = (x^2+y^2)e^(-y^2)

OpenStudy (watchmath):

\(\nabla f =(e^z,\cos x, ee^z)\)

OpenStudy (anonymous):

I'm having trouble getting from DuF(x,y)=Fx(x,y)a + fy(x,y)b

OpenStudy (watchmath):

Sorry \(\nabla f =<e^z,\cos x, xe^z>\)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=grad++%28xe^z+%2B+ycos%28x%29%29 is that the proper form for wolframalpha?